OpenStudy (anonymous):

Find the geometric mean of 20 and 5. A. 10 B. 4 C. 12.5 D. 25 I am in Geometry, but Mean,Median and Mode still haunt me! I thought I got out of this stuff in 2nd grade! haha

OpenStudy (anonymous):

\[\sqrt{20\times 5}=\sqrt{100}=10\]